OpenStudy (anonymous):

Solve for x: (there should be two answers) x(x + 5) = 3(x + 1) Show steps please..

OpenStudy (agent47):

x^2+5x=3x+3 x^2+5x-3x-3=0 x^2+2x-3=0 (x-1)(x-2)=0

OpenStudy (anonymous):

 equate the factors to zero tehn solve for x